OpenStudy (anonymous):

If f(n+1)=f(n) + 1/3 and f(1)=0 what is f(79)? I know this is super easy but I can't think of how to start it.

hartnn (hartnn):

f(n+1) = f(n) +1/3 is equivalent to f(n+1) = f(1) + n*(1/3) do you need to derive this result ?

hartnn (hartnn):

there ?

OpenStudy (anonymous):

Nevermind. Lol that wasn't it anyway. I ended up figuring it out. Solution to my problem: n=1 then f(n)=0, f(2)=f(1)+1/3 f(1)=0 f(2)=1/3 f(3)=2/3 f(4)=1 or 3/3 By this time you can see the pattern and we can see that for f(n)=(n-1)*1/3 So for f(79) we can just plug in and see that its 26.

OpenStudy (anonymous):

Like I said, easy. Just a brainfart. But thanks for trying to help.

hartnn (hartnn):

yeah i got the same equation algebraically, not by noticing the 'pattern'